# Remarks on Dirac–Bergmann algorithm, Dirac’s conjecture and the extended Hamiltonian

### v2

@@ 4.2.2 Generators of spacetime diffeomorphisms
To examine the validity of Dirac’s conjecture within the ADM formalism of GR we need to construct the Castellani generator. As in the Yang–Mills case (33), the part of the generator that contains no derivatives of the gauge parameter differs from the secondary constraint by a specific linear combination of constraints with fixed coefficients:
@@
Such transformations are not compatible with the general covariance of GR. Thus, the Dirac conjecture again holds only in the restricted sense of transformations of states of the gravitational field at some fixed instant of time. To make it work in a usual sense we again need to switch to the extended Hamiltonian.
We have 10 variables in total, 8 constraints, and 4 independent diffeomorphisms $\xi^{\mu}$. As a result, we find $N_{\text{dyn}}=2$, $N_{\text{gauge}}=4$, and $N_{\text{con}}=4$. Six of the ten modes are physical, while four correspond to the arbitrariness in the choice of coordinate system. The two dynamical modes describe the two polarizations of a gravitational wave, while the four constrained modes correspond to the four types of metric perturbations around a background in cosmological perturbation theory: two scalar types and one vector type, the latter having two independent components.
@@ 4.2.3 The extended Hamiltonian approach
Upon introducing the extended Hamiltonian, the first-order action and the corresponding equations of motion again differ from the total Hamiltonian case by a reparameterization of those variables that are second class with respect to the primary constraints. In the present case, these variables are the lapse and shift functions:
$N_{\mu}\equiv N^{\scriptscriptstyle E}_{\mu}+\lambda^{g}_{\mu},\qquad\lambda^{\scriptscriptstyle N}_{\mu}\equiv\lambda^{\scriptscriptstyle E}_{\mu}+\dot{\lambda}^{g}_{\mu}$ (61)
@@
$H_{E}=\int d^{3}\vec{x}\bigl(N^{\scriptscriptstyle E}_{\mu}\>\mathcal{H}^{\mu}+\lambda^{\scriptscriptstyle E}_{\mu}\>\pi^{\mu}+\lambda^{g}_{\mu}\mathcal{H}^{\mu}\bigr).$ (62)
@@
Let us begin with the correct interpretation of the extended Hamiltonian approach. It is, again, equivalent to modifying the functional dependence of the canonical gauge invariants of the theory on the generalized coordinates and velocities. To simplify the analysis, we introduce two types of objects: functionals of the canonical variables expressed in terms of $N^{\scriptscriptstyle E}_{\mu}$ in exactly the same way as they would be expressed in terms of $N_{\mu}$ if the Hamiltonian was total, and the corresponding functionals for the extended Hamiltonian. We will denote objects of the first type with the index T, while objects of the second type will be written without any index. With this interpretation, the standard definitions are no longer satisfied for the spatial momenta. Instead, we have
$\pi^{ij}=\Omega\pi_{\scriptscriptstyle T}^{ij}+\Omega L^{ij},$ (63)
@@
$\Omega=\frac{N^{\scriptscriptstyle E}}{N^{\scriptscriptstyle E}+\lambda^{g}_{0}},\qquad L^{ij}=\frac{\sqrt{\gamma}}{2\varkappa}\bigl(\Lambda^{l}_{l}\>\gamma^{ij}-\Lambda^{ij}\bigr),\qquad\Lambda_{ij}=\frac{1}{2N^{\scriptscriptstyle E}}\bigl(\overset{3}{\nabla}_{i}\lambda^{g}_{j}+\overset{3}{\nabla}_{j}\lambda^{g}_{i}\bigr).$
This immediately leads to modified expressions for the secondary constraints:
$\mathcal{H}_{0}=\Omega^{2}\>\mathcal{H}_{\scriptscriptstyle T}{}_{0}+\Omega^{2}\>\tau_{0},\qquad\qquad\mathcal{H}^{i}=\Omega\>\mathcal{H}_{\scriptscriptstyle T}{}^{i}+\tau^{i},$ (64)
@@
$\tau_{0}=\frac{2\varkappa}{\sqrt{\gamma}}\bigl(\Lambda^{kl}\Lambda_{kl}-\frac{1}{2}\Lambda^{l}{}_{l}{}^{2}\bigr)+\frac{2\varkappa}{\sqrt{\gamma}}\bigl(2\pi_{kl}\Lambda^{kl}+\pi^{l}{}_{l}\Lambda^{k}_{k}\bigr),\qquad\tau^{i}=-2\pi^{ki}\partial_{k}\Omega-2\sqrt{\gamma}\overset{3}{\nabla}_{k}\biggl(\frac{\Omega L^{ki}}{\sqrt{\gamma}}\biggr).$
The quantities $\mathcal{H}_{\mu}$ are the new canonical invariants: $\mathcal{H}_{\mu}=0$, whereas $\mathcal{H}_{\scriptscriptstyle T}{}_{\mu}$ need not vanish on the constraint surface. Accordingly, the correspondence between the Lagrangian and Hamiltonian formalisms is given by (54), and not by
$G^{00}=-\frac{\varkappa}{2\mathfrak{N}^{2}\sqrt{\gamma}}\>\mathcal{H}_{\scriptscriptstyle T}{}_{0},\qquad\qquad G^{i0}=\frac{\varkappa}{2\mathfrak{N}\sqrt{\gamma}}\biggl(\mathcal{H}_{\scriptscriptstyle T}^{i}+\frac{\mathfrak{N}^{i}}{\mathfrak{N}}\mathcal{H}_{\scriptscriptstyle T}^{0}\biggr).\qquad$ (65)
@@
$g^{00}=-\frac{1}{(N^{\scriptscriptstyle E}+\lambda^{g}_{0}){}^{2}},\qquad g^{0k}=\frac{N^{k}+\lambda^{g}{}^{k}}{(N^{\scriptscriptstyle E}+\lambda^{g}_{0})^{2}}.$ (66)
If we count the new Lagrange multipliers $\lambda^{g}_{\mu}$ as part of the total number of modes, then the degree of freedom counting gives $N_{dyn}=2$, $N_{con}=4$, $N_{gauge}=8$. The new Lagrange multipliers $\lambda^{g}_{\mu}$ constitute four additional gauge modes that have been introduced into the theory. They provide extra gauge freedom in the form of (60), which is generated by $G_{D}$. As in the previous cases, these new gauge modes can also be obtained by making the substitution (61) directly in the Lagrangian (47).
In the incorrect approach, one assumes that the components of the Einstein tensor are related to the secondary constraints by the formulas (65). As in the Yang–Mills theory, this interpretation leads to additional arbitrary matter contributions, in the sense that an effective nonzero energy–momentum tensor appears on the right-hand side of the Einstein equations. From (64) we obtain
@@
The spatial EFE’s receive a nonzero contribution as well, but we do not write it out explicitly because the expression is rather cumbersome. The important point is that, in general, all components of this energy–momentum tensor are nonzero. They vanish when $\lambda^{g}_{\mu}=0$, i.e. when the extended Hamiltonian coincides with the total one. The freedom to choose this “matter” arbitrarily is precisely what underlies the additional gauge modes. Once again they “eat up” the constrained modes, and we obtain $N_{dyn}=2$, $N_{con}=0$, $N_{gauge}=8$.
@@ 5 Generalization to all canonical mechanical gauge systems
We claim that the conclusions drawn from the examples above are universal for all canonical gauge theories in which “the gauge hits twice”. We will prove this statement for mechanical systems.
It turns out that the transition from the total Hamiltonian to the extended Hamiltonian can always be achieved by introducing new spurious variables through reparameterization of the original canonical variables in the first–order action. These new variables subsequently appear as Lagrange multipliers for the secondary constraints in the extended Hamiltonian. This procedure is in complete analogy with what occurs in the toy model, Yang–Mills theory, and the ADM formulation of general relativity. Moreover, it must be applied only to those canonical variables that are second-class with respect to the set of primary constraints.
Suppose we consider a system in which all constraints are first–class. The property that “the gauge hits twice” means that to each primary first-class constraint there corresponds a secondary first-class constraint generated by the requirement of its preservation in time. In the total Hamiltonian approach, the most general first–order Lagrangian for a constrained system can be written as
$L^{(1)}=\dot{q}^{a}p_{a}+\dot{q}^{A}p_{A}-H_{can}(q,p^{B})-\lambda^{1}_{a}[p^{a}-f^{a}(q,p^{B})],$ (68)
@@
It is assumed that the system does not have tertiary constraints [^20] We shall assume that the system has no constraints of the $\chi^{n}$ type either.. One may choose the secondary constraints as
$\phi^{a}_{2}\equiv\{p^{a}-f^{a}(q,p^{A}),H_{can}\}=-\frac{\partial H_{can}}{\partial q_{a}}-\frac{\partial f^{a}}{\partial q_{B}}\frac{\partial H_{can}}{\partial p^{B}}+\frac{\partial f^{a}}{\partial p^{i}}\frac{\partial H_{can}}{\partial q^{i}}.$ (69)
The form of the redefinition is motivated by the need for Dirac’s conjecture to hold in the standard interpretation of gauge transformations, as in the examples discussed above. To this end, we must introduce an additional gauge freedom into the dynamics of those canonical variables that are second class with respect to the primary constraints. This, in turn, ensures an additional independent contribution of primary constraints to the Castellani generator (15), giving rise to transformations of the form [^21] Note that if a variable $z$ commutes with $\phi_{1}^{a}$, then the functions $f^{a}$ do not depend on its conjugate variable, and the corresponding derivative in (70) vanishes, ($\frac{\partial f^{a}}{\partial p_{z}}=0$).
$\delta q_{a}=\alpha_{b}\{q_{a},\phi_{1}^{b}\}=\alpha^{a},\quad\delta q_{A}=\alpha_{b}\{q_{A},\phi_{1}^{b}\}=-\alpha_{b}\>\frac{\partial f^{b}}{\partial p^{A}},\quad\delta p^{i}=\alpha_{b}\{p^{i},\phi_{1}^{b}\}=\alpha_{b}\>\frac{\partial f^{b}}{\partial q_{i}}.$ (70)
@@
$q_{i}=q^{\scriptscriptstyle E}_{i}-\lambda^{2}_{a}\{q_{i},\phi_{1}^{a}\},\qquad p^{i}=p_{\scriptscriptstyle E}^{i}-\lambda^{2}_{a}\{p^{i},\phi_{1}^{a}\},$ (71)
@@
If we perform the substitution (71) in the first-order Lagrangian, expand in a series in $\lambda^{2}_{a}$, and keep only the terms up to first order neglecting the total derivative term, we obtain
@@
which is the first-order Lagrangian of the extended Hamiltonian system. All the contributions that are proportional to the primary constraints are absorbed into the Lagrangian multipliers $\lambda_{a}^{\scriptscriptstyle E}{}^{1}$.
This observation proves that all our arguments apply to any system in which the “gauge hits twice”. Under the correct interpretation of the Extended Hamiltonian the number of physical modes is preserved, but we acquire $K$ new pure gauge modes $\lambda^{2}_{a}$.
This statement becomes essentially trivial when the functions $f^{a}$ depend solely on the coordinates $q^{a}$. In this case, the functions $f^{a}$ commute with the canonical Hamiltonian, and the secondary constraints take the form $\phi^{a}_{2}=-\frac{\partial H_{can}}{\partial q^{a}}(q,p^{A})$, so that they do not depend on the constrained momenta $p_{a}$. Requiring the constraints $\phi^{a}_{2}$ to be first-class then implies that they are also independent of the unresolved coordinates $q_{a}$. Consequently, the secondary constraints must enter the canonical Hamiltonian linearly, multiplied by the generalized coordinates that are second class with respect to them, namely,
@@
This structure is precisely what is observed in the toy model (9), in Yang–Mills theory (30), and in the ADM formalism (48). Therefore, the substitution (71) amounts to a shift $q_{a}\,\phi^{a}_{2}(q_{B},p^{A})\rightarrow-(q_{a}+\lambda_{a}^{2})\,\phi^{a}_{2}(q_{B},p^{A})$.
@@
If we express $\phi^{a}_{2}$ in terms of the generalized velocities, we obtain Lagrangian constraints, i.e. the Euler–Lagrange equations for the variables $q_{a}$, since $\dot{p}^{a}=-\frac{\partial H_{T}}{\partial q_{a}}=\frac{\partial L}{\partial q_{a}},\ \dot{f}^{a}=\frac{d}{dt}\frac{\partial L}{\partial\dot{q}_{a}}$. With the incorrect interpretation of the Extended Hamiltonian, however, the newly introduced $K$ pure gauge modes take over the role of the original constrained ones, while the Lagrangian constraints get modified. This happens because the functions $\phi^{a}_{2}$ are then interpreted as functions of $q^{\scriptscriptstyle E}$ and $\dot{q}^{\scriptscriptstyle E}$: $\phi^{2}_{a}=\phi^{2}_{a}(q^{\scriptscriptstyle E},\dot{q}^{\scriptscriptstyle E})$, rather than of the original variables $\phi^{2}_{a}=\phi^{2}_{a}(q,\dot{q})$. As a result, the right–hand sides of the corresponding Euler–Lagrange equations acquire contributions that depend on the new Lagrange multipliers:
@@
$\frac{\delta S}{\delta q_{a}}\biggl|_{q=q^{\scriptscriptstyle E}}=\phi^{a}_{2}(q^{\scriptscriptstyle E},\dot{q}^{\scriptscriptstyle E})=-\frac{\partial\phi^{a}_{2}}{\partial q_{i}}\,\lambda^{2}_{a}\{q^{\scriptscriptstyle E}_{i},\phi^{a}_{1}\}\Big|_{p_{i}=\frac{\partial L}{\partial\dot{q}_{i}}}-\frac{\partial\phi^{a}_{2}}{\partial\dot{q}_{i}}\,\frac{d}{dt}\!\left(\lambda^{2}_{a}\{q^{\scriptscriptstyle E}_{i},\phi^{a}_{1}\}\Big|_{p_{i}=\frac{\partial L}{\partial\dot{q}_{i}}}\right)\neq 0.$ (73)

### v3

@@ 4.2.2 Generators of spacetime diffeomorphisms
To examine the validity of Dirac’s conjecture within the ADM formalism of GR we need to construct the Castellani generator. As in the Yang-Mills case (33), the part of the generator that contains no derivatives of the gauge parameter differs from the secondary constraint by a specific linear combination of constraints with fixed coefficients:
@@
Such transformations are not compatible with the general covariance of GR. Thus, the Dirac conjecture again holds only for transformations of the gravitational field’s state at a fixed instant. To make it work in a usual sense, we again need to switch to the extended Hamiltonian.
We have 10 variables in total, 8 constraints, and 4 independent diffeomorphisms $\xi^{\mu}$. As a result, we find $N_{\text{dyn}}=2$, $N_{\text{gauge}}=4$, and $N_{\text{con}}=4$. Six of the ten modes are physical, while four correspond to the arbitrariness in the choice of the coordinate system. The two dynamical modes describe the two polarizations of a gravitational wave, while the four constrained modes correspond to the four types of metric perturbations around a background in cosmological perturbation theory: two scalar types and one vector type, the latter having two independent components.
@@ 4.2.3 The extended Hamiltonian approach
Upon introducing the extended Hamiltonian, the first-order action and the corresponding equations of motion differ from the total Hamiltonian case by a reparameterization of those variables that are second-class with respect to the primary constraints. In the present case, these variables are the lapse and shift functions:
$N_{\mu}\equiv N^{\scriptscriptstyle E}_{\mu}+\lambda^{\scriptscriptstyle\mathcal{H}}_{\mu},\qquad\lambda^{\scriptscriptstyle N}_{\mu}\equiv\lambda^{\scriptscriptstyle E}_{\mu}+\dot{\lambda}^{\scriptscriptstyle\mathcal{H}}_{\mu}$ (61)
@@
$H_{E}=\int d^{3}\vec{x}\bigl(N^{\scriptscriptstyle E}_{\mu}\>\mathcal{H}^{\mu}+\lambda^{\scriptscriptstyle E}_{\mu}\>\pi^{\mu}+\lambda^{\scriptscriptstyle\mathcal{H}}_{\mu}\mathcal{H}^{\mu}\bigr).$ (62)
@@
Let us begin with the correct interpretation of the extended Hamiltonian approach. All objects in the Hamiltonian formulation can be expressed as functionals of the Lagrangian variables – the spatial metric, lapse, and shifts. The extended Hamiltonian formulation is, once again, equivalent to modifying the functional dependence of the canonical gauge invariants of the theory on the generalized coordinates and velocities. To simplify the analysis, we introduce two types of objects: functionals of the Lagrangian variables arising in the total Hamiltonian formulation and the corresponding functionals evaluated at the extended Hamiltonian’s variables $N_{\mu}^{\scriptscriptstyle E}$. We denote objects of the second type with the index $E$, while those of the first type are written without it. For example,
$K_{ij}=\frac{1}{2N}\biggl(\overset{3}{\nabla}_{i}N_{j}+\overset{3}{\nabla}_{j}N_{i}-\dot{\gamma}_{ij}\biggr)\quad\text{and}\quad K^{\scriptscriptstyle E}_{ij}=\frac{1}{2N^{\scriptscriptstyle E}}\biggl(\overset{3}{\nabla}_{i}N^{\scriptscriptstyle E}_{j}+\overset{3}{\nabla}_{j}N^{\scriptscriptstyle E}_{i}-\dot{\gamma}_{ij}\biggr).$ (63)
The difference is due to the substitution (61). Within the correct approach, the standard definitions are no longer satisfied for the spatial momenta. Instead, we have
$\pi^{ij}=\Omega\pi_{\scriptscriptstyle E}^{ij}+\Omega L^{ij},$ (64)
@@
$\Omega=\frac{N^{\scriptscriptstyle E}}{N^{\scriptscriptstyle E}+\lambda^{\scriptscriptstyle\mathcal{H}}_{0}},\qquad L^{ij}=\frac{\sqrt{\gamma}}{2\varkappa}\bigl(\Lambda^{l}_{l}\>\gamma^{ij}-\Lambda^{ij}\bigr),\qquad\Lambda_{ij}=\frac{1}{2N^{\scriptscriptstyle E}}\bigl(\overset{3}{\nabla}_{i}\lambda^{\scriptscriptstyle\mathcal{H}}_{j}+\overset{3}{\nabla}_{j}\lambda^{\scriptscriptstyle\mathcal{H}}_{i}\bigr).$
This immediately leads to modified expressions for the Lagrangian constraints:
$\mathcal{H}_{0}=\Omega^{2}\>\mathcal{H}^{\scriptscriptstyle E}{}_{0}+\Omega^{2}\>\tau_{0},\qquad\qquad\mathcal{H}^{i}=\Omega\>\mathcal{H}_{\scriptscriptstyle E}{}^{i}+\tau^{i},$ (65)
@@
$\tau_{0}=\frac{2\varkappa}{\sqrt{\gamma}}\bigl(\Lambda^{kl}\Lambda_{kl}-\frac{1}{2}\Lambda^{l}{}_{l}{}^{2}\bigr)+\frac{2\varkappa}{\sqrt{\gamma}}\bigl(2\pi^{\scriptscriptstyle E}_{kl}\Lambda^{kl}+\pi_{\scriptscriptstyle E}^{l}{}_{l}\Lambda^{k}_{k}\bigr),\qquad\tau^{i}=-2\pi_{\scriptscriptstyle E}^{ki}\partial_{k}\Omega-2\sqrt{\gamma}\overset{3}{\nabla}_{k}\biggl(\frac{\Omega L^{ki}}{\sqrt{\gamma}}\biggr).$
The quantities $\mathcal{H}_{\mu}$ serve as the canonical invariants in this formulation and satisfy $\mathcal{H}_{\mu}=0$, whereas $\mathcal{H}^{\scriptscriptstyle E}{}_{\mu}$ need not vanish on the constraint surface. Accordingly, the correspondence between the Lagrangian and Hamiltonian formalisms is given by (54), rather than by
$G^{00}=-\frac{\varkappa}{2N^{\scriptscriptstyle E}{}^{2}\sqrt{\gamma}}\>\mathcal{H}^{\scriptscriptstyle E}{}_{0},\qquad\qquad G^{i0}=\frac{\varkappa}{2N^{\scriptscriptstyle E}\sqrt{\gamma}}\biggl(\mathcal{H}_{\scriptscriptstyle E}^{i}+\frac{N_{\scriptscriptstyle E}{}^{i}}{N^{\scriptscriptstyle E}}\mathcal{H}_{\scriptscriptstyle E}^{0}\biggr).\qquad$ (66)
@@
$g^{00}=-\frac{1}{(N^{\scriptscriptstyle E}+\lambda^{\scriptscriptstyle\mathcal{H}}_{0}){}^{2}},\qquad g^{0k}=\frac{N^{k}+\lambda^{\scriptscriptstyle\mathcal{H}}{}^{k}}{(N^{\scriptscriptstyle E}+\lambda^{\scriptscriptstyle\mathcal{H}}_{0})^{2}}.$ (67)
If we count the new Lagrange multipliers $\lambda^{\scriptscriptstyle\mathcal{H}}_{\mu}$ as part of the total number of modes, then the degree of freedom counting gives $N_{dyn}=2$, $N_{con}=4$, $N_{gauge}=8$. The new Lagrange multipliers $\lambda^{\scriptscriptstyle\mathcal{H}}_{\mu}$ constitute four additional gauge modes that have been introduced into the theory. They provide extra gauge freedom in the form of (60), which is generated by $G_{D}$. As in the previous cases, these new gauge modes can also be obtained by making the substitution (61) directly in the Lagrangian (47).
In the incorrect approach, one assumes that the components of the Einstein tensor are related to the secondary constraints by the formulas (66). As in the Yang-Mills theory, this interpretation leads to additional arbitrary matter contributions, in the sense that an effective nonzero energy–momentum tensor appears on the right-hand side of the Einstein equations. From (65) we obtain
@@
The spatial Einstein equations also receive a nonvanishing contribution as well, but we do not write it out explicitly since the resulting expression is rather cumbersome. The important point is that, in general, all components of this energy-momentum tensor are nonzero. They vanish when $\lambda^{\scriptscriptstyle\mathcal{H}}_{\mu}=0$, i.e., when the extended Hamiltonian coincides with the total one. The freedom to choose this “matter” arbitrarily is precisely what underlies the additional gauge modes. Once again, they make the constrained modes unphysical, and we obtain $N_{dyn}=2$, $N_{con}=0$, $N_{gauge}=8$.
@@ 5 Generalization to all canonical mechanical gauge systems
We claim that the conclusions drawn from the examples above are universal for all canonical gauge theories in which “the gauge hits twice”. We will prove this statement for mechanical systems. It turns out that the transition from the total Hamiltonian to the extended Hamiltonian can always be achieved by introducing new spurious variables through reparameterization of the original canonical variables in the first-order action. These new variables subsequently appear as Lagrange multipliers for the secondary constraints in the extended Hamiltonian. This procedure is in complete analogy with what occurs in the toy model, Yang-Mills theory, and the ADM formulation of general relativity. Moreover, it must be applied only to those canonical variables that are second-class with respect to the set of primary constraints.
Suppose we consider a system in which all constraints are first-class. The property that “the gauge hits twice” means that to each primary first-class constraint, there corresponds a secondary first-class constraint generated by the requirement of its preservation in time. It is assumed that the system does not have tertiary constraints [^20] We shall assume that the system has no constraints of the $\chi^{n}$ type either.. As mentioned before, the equations of motion for our system follow from the first-order Lagrangian of a constrained system, which has the form
$L^{(1)}=\dot{q}_{a}p^{a}+\dot{q}_{A}p^{A}-H_{can}(q,p^{B})-\lambda^{1}_{a}[p^{a}-f^{a}(q,p^{B})],$ (69)
@@
One may choose the secondary constraints as
$\phi^{a}_{2}(q,p^{A})\equiv\{p^{a}-f^{a}(q,p^{A}),H_{can}(q,p^{A})\}=-\frac{\partial H_{can}}{\partial q_{a}}-\frac{\partial f^{a}}{\partial q_{B}}\frac{\partial H_{can}}{\partial p^{B}}+\frac{\partial f^{a}}{\partial p^{i}}\frac{\partial H_{can}}{\partial q^{i}}.$ (70)
The form of the reparameterization is motivated by the need for Dirac’s conjecture to hold in the standard interpretation of gauge transformations, as in the examples discussed above. To this end, we must introduce an additional gauge freedom into the dynamics of those canonical variables that are second-class with respect to the primary constraints. This, in turn, ensures an additional independent contribution of primary constraints to the Castellani generator (15), which gives rise to transformations [^21] Note that if a variable $z$ commutes with $\phi_{1}^{a}$, then the functions $f^{a}$ do not depend on its conjugate variable, and the corresponding derivative in (71) vanishes.
$\delta q_{a}=\alpha_{b}\{q_{a},\phi_{1}^{b}\}=\alpha_{a},\quad\delta q_{A}=\alpha_{b}\{q_{A},\phi_{1}^{b}\}=-\alpha_{b}\>\frac{\partial f^{b}}{\partial p^{A}},\quad\delta p^{i}=\alpha_{b}\{p^{i},\phi_{1}^{b}\}=\alpha_{b}\>\frac{\partial f^{b}}{\partial q_{i}}.$ (71)
@@
$q_{i}=q^{\scriptscriptstyle E}_{i}-\lambda^{2}_{a}\{q^{\scriptscriptstyle E}_{i},\phi_{1}^{a}\},\qquad p^{i}=p_{\scriptscriptstyle E}^{i}-\lambda^{2}_{a}\{p_{\scriptscriptstyle E}^{i},\phi_{1}^{a}\},$ (72)
@@
By performing the substitution (72) into the first-order Lagrangian and keeping terms only up to first order in $\lambda^{2}_{a}$, while omitting total derivative terms, we obtain:
@@
which is the first-order Lagrangian of the extended Hamiltonian system. All the contributions that are proportional to the primary constraints are absorbed into the term $\lambda_{a}^{\scriptscriptstyle E}{}^{1}\phi^{a}_{1}$. This observation proves that all our arguments apply to any system in which ”the gauge hits twice”. Under the correct interpretation of the Extended Hamiltonian the number of physical modes is preserved, but we acquire $K$ new pure gauge modes $\lambda^{2}_{a}$.
This statement becomes essentially trivial when the functions $f^{a}$ depend solely on the coordinates $q_{a}$. In this case, the functions $f^{a}$ commute with the canonical Hamiltonian, and the secondary constraints take the form $\phi^{a}_{2}=-\frac{\partial H_{can}}{\partial q^{a}}(q,p^{A})$. Requiring the constraints $\phi^{a}_{2}$ to be first-class implies that they are independent of the unresolved coordinates $q_{a}$, as they do not depend on the constrained momenta $p_{a}$ by construction (70). Consequently, the secondary constraints must enter the canonical Hamiltonian linearly, multiplied by the generalized coordinates that are second-class with respect to them, namely,
@@
This structure is precisely what is observed in the toy model (9), in Yang-Mills theory (30), and in the ADM formalism (48). Therefore, the substitution (72) amounts to a shift $q_{a}\,\phi^{a}_{2}\rightarrow-(q^{\scriptscriptstyle E}_{a}+\lambda_{a}^{2})\,\phi^{a}_{2}$.
If we express $\phi^{a}_{2}$ in terms of the generalized velocities, we obtain Lagrangian constraints, i.e., the Euler–Lagrange equations for the variables $q_{a}$; since $\dot{p}^{a}=-\frac{\partial H_{T}}{\partial q_{a}}=\frac{\partial L}{\partial q_{a}}$ and $\dot{f}^{a}=\frac{d}{dt}\frac{\partial L}{\partial\dot{q}_{a}}$, the following relation holds
$\frac{\delta S}{\delta q_{a}}=\phi^{a}_{2}(q,f^{B}(q,\dot{q}_{A},\dot{q}_{b})).$ (74)
@@
With the incorrect interpretation of the Extended Hamiltonian, however, the newly introduced $K$ pure gauge modes take over the role of the original constrained ones, while the Lagrangian constraints are modified. This occurs because, instead of identifying the Lagrangian constraints with the original functions $\phi^{a}_{2}(q,f^{B})$, we interpret these constraints as the same functions evaluated at the extended Hamiltonian variables $q^{\scriptscriptstyle E}$ and $\dot{q}^{\scriptscriptstyle E}$: $\phi^{a}_{2}=\phi^{a}_{2}(q^{\scriptscriptstyle E},f^{B}(q,\dot{q}^{\scriptscriptstyle E}_{A},\dot{q}^{\scriptscriptstyle E}_{b}))$. The right-hand sides of the corresponding Euler–Lagrange equations become spurious, as they now depend on the new Lagrange multipliers:
@@
$\frac{\delta S}{\delta q_{a}}\biggl|_{q=q^{\scriptscriptstyle E}}=\phi^{a}_{2}(q^{\scriptscriptstyle E},f^{B}(q,\dot{q}^{\scriptscriptstyle E}_{A},\dot{q}^{\scriptscriptstyle E}_{b}))=-\frac{\partial\phi^{a}_{2}}{\partial q_{i}}\,\lambda^{2}_{a}\{q^{\scriptscriptstyle E}_{i},\phi^{a}_{1}\}\Big|_{p^{i}=\frac{\partial L}{\partial\dot{q}_{i}}}-\frac{\partial\phi^{a}_{2}}{\partial\dot{q}_{i}}\,\frac{d}{dt}\!\left(\lambda^{2}_{a}\{q^{\scriptscriptstyle E}_{i},\phi^{a}_{1}\}\Big|_{p^{i}=\frac{\partial L}{\partial\dot{q}_{i}}}\right)\neq 0.$ (75)